A man loans invests PhP 2,000 at a certain rate and PhP 3,000 at a rate of 2/3 as high. His annual income from the two investments is PhP 120. Find the interest rates.
Answer by richwmiller(17219)

2000x+3000*2/3x=120
2000x+2000x=120
4000x=120
100x=3
x=3/100 =3%
2/3x=2/100=2%